Find the inverse of the following matrix using row reduction method

[ 1 -1 2]
[ 3 2 4]
[ 0 1 -2]
1
Expert's answer
2021-07-19T17:21:56-0400


To find the inverse matrix, augment it with the identity matrix and perform row operations trying to make the identity matrix to the left. Then to the right will be the inverse matrix.



So, augment the matrix with the identity matrix:



"\begin{bmatrix}
 1& -1 &2 & \kern{3 em} 1&0&0\\
 3& 2 &4 & \kern{3 em} 0&1&0\\
 0& 1 &-2 & \kern{3 em} 0&0&1
\end{bmatrix}"




"\rightarrow" "R_2 = R_2 - 3R_1"


"\begin{bmatrix}
 1& -1 &2 & \kern{3 em} 1&0&0\\
 0& 5 &-2 & \kern{3 em} -3&1&0\\
 0& 1 &-2 & \kern{3 em} 0&0&1
\end{bmatrix}"





"\rightarrow" "R_2 = \dfrac{R_2 }{5}"


"\begin{bmatrix}
 1& -1 &2 & \kern{3 em} 1&0&0\\
 0& 1 &-2/5 & \kern{3 em} -3/5&1/5&0\\
 0& 1 &-2 & \kern{3 em} 0&0&1
\end{bmatrix}"






"\rightarrow" "R_1 = R_1 +R_2"


"\begin{bmatrix}
 1& 0 &8/5 & \kern{3 em} 2/5 &1/5&0\\
 0& 1 &-2/5 & \kern{3 em} -3/5&1/5&0\\
 0& 1 &-2 & \kern{3 em} 0&0&1
\end{bmatrix}"






"\rightarrow" "R_3 = R_3 - R_2"


"\begin{bmatrix}
 1& 0 &8/5 & \kern{3 em} 2/5&1/5&0\\
 0& 1 &-2/5 & \kern{3 em} -3/5&1/5&0\\
 0& 0 &-8/5 & \kern{3 em} 3/5&-1/5&1
\end{bmatrix}"






"\rightarrow" "R_3 = -\dfrac{5}{ 8}R_3"


"\begin{bmatrix}
 1& 0 &8/5 & \kern{3 em} 2/5&1/5&0\\
 0& 1 &-2/5 & \kern{3 em} -3/5&1/5&0\\
 0& 0 &1 & \kern{3 em} -3/8&1/8&-5/8
\end{bmatrix}"






"\rightarrow" "R_1 = R_1 - \dfrac{8}{5}R_3"


"\begin{bmatrix}
 1& 0 &0 & \kern{3 em} 1&0&1\\
 0& 1 &-2/5 & \kern{3 em} -3/5&1/5&0\\
 0& 0 &1 & \kern{3 em} -3/8&1/8&-5/8
\end{bmatrix}"







"\rightarrow" "R_2 = R_2 + \dfrac{2}{5}R_3"


"\begin{bmatrix}
 1& 0 &0 & \kern{3 em} 1&0&1\\
 0& 1 &0 & \kern{3 em} -3/4&1/4&-1/4\\
 0& 0 &1 & \kern{3 em} -3/8&1/8&-5/8
\end{bmatrix}"




On the left is the identity matrix. On the right is the inverse matrix.


The inverse matrix is = "\begin{bmatrix}
 1&0&1\\
 -3/4&1/4&-1/4\\
 -3/8&1/8&-5/8
\end{bmatrix}"